What are the solutions to this system of equations?
x^2 + y^2 = 5
y = 2x

Answer by rothauserc(4718)

substitute 2x for y in the first equation
x^2 + (2x)^2 = 5
x^2 + 4x^2 = 5
5x^2 = 5
x^2 = 1
x = square root(1) = 1 or -1
y = 2x = 2 or -2
we have two solution sets
x = +1, y = +2 or
x = -1, y = -2